OpenStudy (anonymous):

how do you evaluate the indefinite integral of Sin Pi(t) dt using the substitution rule?

OpenStudy (anonymous):

i was thinking U=sin Pi(t) and du= -cos Pi(t) but i dont know if that right

OpenStudy (espex):

If your problem is \[\int{} \pi \sin{t}\space dt\] then you should be able to just integrate directly.

OpenStudy (anonymous):

no its \[\int\limits_{?}^{?} \sin pi (t) dt\]

OpenStudy (anonymous):

i just got it. thanks anyways!